OpenStudy (anonymous):

A mass m = 48 g is suspended from a spring with spring constant k = 4.0 N/m. 1. Determine the force acting on the mass when it is displaced 7 cm from its equilibrium position. If the mass is then released, what is the period of oscillation? What is the frequency? 2. Mary observes damped oscillations of a glider on an air track. She observed that the amplitude decreased to 30% of its original value after 12 seconds. What is the decay constant for the motion of the glider?

OpenStudy (anonymous):

Use Hooke's Law, F = kx F is force, k is spring constant, x is displacement. Careful with units on this one, using Newtons means kg and m are needed for mass and displacement..

OpenStudy (anonymous):

and for time period, \[T = 2 \pi \sqrt{\frac{ m }{ k }}\]

OpenStudy (anonymous):

And of course \[f = \frac{ 1 }{ T }\]

OpenStudy (anonymous):

reported abusive

OpenStudy (anonymous):

their username..?!

OpenStudy (anonymous):

thanks! :D